# Internal Causal Mechanisms Robustly Predict Language Model Out-of-Distribution Behaviors

### v1

@@ 1 Introduction
As language models are increasingly deployed in high-stakes, open-ended scenarios, it is critical that we have a handle on how to prevent incorrect or unsafe outputs across a wide range of unanticipated user inputs. However, behavioral testing usually cannot be comprehensive given the compositionally large input space of a natural language task and the potential for distribution shifts across deployed environments. The traditional approach is to estimate the correctness of model predictions using confidence scores (Hendrycks & Gimpel, 2017; Lakshminarayanan et al., 2017; Guo et al., 2017). A more recent line of work uses internal representations of the model (usually supervised probes or use other statistical measures of model activations) to predict the model’s correctness on factual retrieval tasks (Azaria & Mitchell, 2023; Zou et al., 2023; Orgad et al., 2024; Marks & Tegmark, 2024; Gottesman & Geva, 2024; Ferrando et al., 2024; Ji et al., 2024; Yin et al., 2024; Chuang et al., 2024).
@@ 2.2 Correctness estimation
Correctness estimation is a fundamental problem in machine learning evaluation and reliability assessment. It is the problem of predicting whether a model output matches the ground truth. For a classifier, a natural choice is to use probabilistic predictions, i.e., the confidence scores (Hendrycks & Gimpel, 2017; Lakshminarayanan et al., 2017). However, LMs’ outputs are structured predictions which add complexities to calibration (Kuleshov & Liang, 2015), and deep neural models are only well-calibrated to in-domain inputs after temperature scaling (Guo et al., 2017; Desai & Durrett, 2020; Kadavath et al., 2022; OpenAI, 2024; Wang et al., 2024). Alternatives have been proposed, such as using verbalized confidence scores for instruction-tuned models (Kadavath et al., 2022; Tian et al., 2023). Most related to ours is the line of work using internal representations to estimate the knowledge stored in LMs (Gottesman & Geva, 2024; Marks & Tegmark, 2024; Ferrando et al., 2024), assess the risk of hallucination (Azaria & Mitchell, 2023; Zou et al., 2023; Du et al., 2024; Chuang et al., 2024; Ji et al., 2024; Yin et al., 2024; Orgad et al., 2024), and predict jailbreaking behaviors (Arditi et al., 2024). However, identifying robust features from internal representations that generalize across tasks and datasets remains a challenge (Orgad et al., 2024). This is the problem that we focus on in this work.
@@ 5 Experiments
Confidence Score ×$\times N$
[…]
Counterfactual Simulation Localization Only ×$\times NK$
[…]
Table 1: A comparison of methods and their requirements. Requires training: whether the method requires learning additional parameters; requires wrong samples: whether supervision from incorrect examples is needed; requires counterfactuals: whether supervision from counterfactual examples is needed; requires decoding: whether decoding at inference time is needed; inference cost: expected number of forward passes needed, with $N$ decoding steps and $K$ counterfactual samples. Localization only indicates the condition is required only for localizing causal variables, but not for correctness prediction; Maybe indicates methods that can apply to both prompt tokens and decoded tokens.
@@
Confidence Score (First N Token) $T$=1 $0.793_{\hskip 1.0pt\pm 0.011}$ $0.935_{\hskip 1.0pt\pm 0.004}$ $0.796_{\hskip 1.0pt\pm 0.023}$ $\mathbf{0.810}_{\hskip 1.0pt\pm 0.009}$ $0.682_{\hskip 1.0pt\pm 0.016}$
$T$=2 $0.804_{\hskip 1.0pt\pm 0.010}$ $0.900_{\hskip 1.0pt\pm 0.016}$ $0.836_{\hskip 1.0pt\pm 0.011}$ $0.795_{\hskip 1.0pt\pm 0.002}$ $0.668_{\hskip 1.0pt\pm 0.023}$
Confidence Score (Answer Token) $T$=1 $0.498_{\hskip 1.0pt\pm 0.002}$ $0.928_{\hskip 1.0pt\pm 0.006}$ $0.795_{\hskip 1.0pt\pm 0.018}$ $\mathbf{0.810}_{\hskip 1.0pt\pm 0.009}$ $0.664_{\hskip 1.0pt\pm 0.015}$
$T$=2 $0.498_{\hskip 1.0pt\pm 0.002}$ $0.870_{\hskip 1.0pt\pm 0.016}$ $0.679_{\hskip 1.0pt\pm 0.015}$ $0.795_{\hskip 1.0pt\pm 0.002}$ $0.657_{\hskip 1.0pt\pm 0.016}$
@@
Table 2: Results of the in-distribution setting. Overall, methods using causal features are the best. Methods using features from the prompt last token might also have high AUC-ROC if it coincides with the location of causal variables. Numbers are reported in mean±$\pm$±std. For each task, the highest AUC-ROC value is in bold. Values that are within a standard deviation are underlined.
@@
      (First N Token) $T$=1 $0.631_{\hskip 1.0pt\pm 0.020}$ $0.777_{\hskip 1.0pt\pm 0.009}$ $0.767_{\hskip 1.0pt\pm 0.012}$ $0.650_{\hskip 1.0pt\pm 0.002}$ $0.874_{\hskip 1.0pt\pm 0.012}$ $0.693_{\hskip 1.0pt\pm 0.012}$ $0.707_{\hskip 1.0pt\pm 0.021}$ $0.770_{\hskip 1.0pt\pm 0.028}$ $0.626_{\hskip 1.0pt\pm 0.000}$ $\underline{0.739}_{\hskip 1.0pt\pm 0.000}$
$T$=2 $0.498_{\hskip 1.0pt\pm 0.003}$ $0.714_{\hskip 1.0pt\pm 0.011}$ $0.752_{\hskip 1.0pt\pm 0.005}$ $0.558_{\hskip 1.0pt\pm 0.009}$ $0.873_{\hskip 1.0pt\pm 0.013}$ $0.651_{\hskip 1.0pt\pm 0.013}$ $\underline{0.761}_{\hskip 1.0pt\pm 0.017}$ $0.679_{\hskip 1.0pt\pm 0.009}$ $0.592_{\hskip 1.0pt\pm 0.000}$ $\underline{0.730}_{\hskip 1.0pt\pm 0.000}$
      (Answer Token) $T$=1 $0.544_{\hskip 1.0pt\pm 0.012}$ $0.771_{\hskip 1.0pt\pm 0.020}$ $0.709_{\hskip 1.0pt\pm 0.017}$ $0.596_{\hskip 1.0pt\pm 0.007}$ $\underline{0.908}_{\hskip 1.0pt\pm 0.016}$ $0.739_{\hskip 1.0pt\pm 0.011}$ $0.707_{\hskip 1.0pt\pm 0.021}$ $0.770_{\hskip 1.0pt\pm 0.028}$ $\mathbf{0.931}_{\hskip 1.0pt\pm 0.018}$ $0.491_{\hskip 1.0pt\pm 0.018}$
$T$=2 $0.498_{\hskip 1.0pt\pm 0.003}$ $0.729_{\hskip 1.0pt\pm 0.016}$ $0.671_{\hskip 1.0pt\pm 0.024}$ $0.521_{\hskip 1.0pt\pm 0.004}$ $\underline{0.906}_{\hskip 1.0pt\pm 0.014}$ $0.504_{\hskip 1.0pt\pm 0.014}$ $\underline{0.761}_{\hskip 1.0pt\pm 0.017}$ $0.679_{\hskip 1.0pt\pm 0.009}$ $\underline{0.902}_{\hskip 1.0pt\pm 0.018}$ $0.484_{\hskip 1.0pt\pm 0.016}$
@@
      (Background Variables) $0.556_{\hskip 1.0pt\pm 0.003}$ $0.597_{\hskip 1.0pt\pm 0.008}$ $0.541_{\hskip 1.0pt\pm 0.006}$ $0.749_{\hskip 1.0pt\pm 0.017}$ $0.812_{\hskip 1.0pt\pm 0.057}$ $0.837_{\hskip 1.0pt\pm 0.007}$ $0.729_{\hskip 1.0pt\pm 0.013}$ $0.698_{\hskip 1.0pt\pm 0.020}$ $0.656_{\hskip 1.0pt\pm 0.009}$ $0.660_{\hskip 1.0pt\pm 0.009}$
@@ 5.1 Setup
We construct a test set with an equal number of correctly and wrongly predicted inputs, which turns the correctness prediction task into a balanced binary classification task. Following the evaluation setup of similar truthfulness prediction tasks (Orgad et al., 2024; Yin et al., 2024), we measure the prediction accuracy using AUC-ROC.
@@ 7 Conclusions
Impact Statements
Limitations
Generalizability to non-linear representations
Requirement of Prior Knowledge of the Task
Indeed, many real-world applications already exhibit shared structures readily exploitable by our methods. Two practical domains exemplify this. The first is LLM evaluation: constructing benchmarks typically requires costly human annotations; thus, a reliable correctness estimator helps benchmark creators prioritize uncertain or error-prone examples. Meanwhile, systematic evaluation often requires structured input—for example, SQuAD adversarial (Jia & Liang, 2017) or GSM-symbolic (Mirzadeh et al., 2024)—which inherently reflects high-level task structures readily exploitable by our framework. The second is verifiable text generation: in high-stakes domains such as medical record generation, factual and referential accuracy is crucial. One strategy is having LLMs generate symbolic references to structured source data that are easy to verify (Hennigen et al., 2024). Such symbolic references act as templated prompts convertible into high-level models, enabling our methods to estimate the correctness of texts produced by these decoding algorithms.
@@ A.1 Task Prompts
Table LABEL:tab:task_prompts summarizes the templates and variables used to construct prompts. Each prompt consists of an instruction, a chat template (for chat models), and an optional prefix attached to model outputs. Prompt design adheres to the following criteria:
@@
We select prompts that achieve reasonably high accuracy, but not perfect accuracy, as discussed in Section 3. This is meant to ensure that (1) the model uses some causal mechanisms to systematically solve the task rather than relying on random guesses, and (2) the model makes some errors, enabling meaningful analysis of conditions under which the model makes correct or incorrect predictions (see task accuracy in Table 3). This balance is especially critical for simpler tasks like IOI, where today’s LLMs like Llama-3-8B-Instruct frequently achieve near-perfect performance.
@@
{upper_bound}>>>{lower_bound} and that they are distinct from {query}.
@@
{CJK} UTF8min [{"市": "{city_a_in_japan}", "国": "Japan", {"市": "{city_b_not_in_japan}",
@@ B.2 Training setup and Hyperparameters
IOI PriceTag RAVEL MMLU UnlearnHP
# of Intervention Dimensions 256 1 1024 4 4
Intervention Localizations Layers
9-12 (position variable)
@@
13-14 (country variable) 17 (position variable) 18 (refusal variable)
@@
For counterfactual simulation, we use 10K pairs randomly sampled from 1024x1024 correct examples as the training data. We use the AdamW optimizer with constant LR = 1e-4 with no weight decay, trained for one epoch.
@@ C.1 Model Variations
     $T$=1 $0.867_{\hskip 1.0pt\pm 0.019}$ $0.638_{\hskip 1.0pt\pm 0.036}$ $0.800_{\hskip 1.0pt\pm 0.016}$
     $T$=2 $0.861_{\hskip 1.0pt\pm 0.023}$ $0.649_{\hskip 1.0pt\pm 0.020}$ $0.898_{\hskip 1.0pt\pm 0.013}$
@@
     $T$=1 $0.853_{\hskip 1.0pt\pm 0.022}$ $0.593_{\hskip 1.0pt\pm 0.012}$ $0.766_{\hskip 1.0pt\pm 0.016}$
     $T$=2 $0.724_{\hskip 1.0pt\pm 0.035}$ $0.607_{\hskip 1.0pt\pm 0.014}$ $0.769_{\hskip 1.0pt\pm 0.019}$
@@
Table 6 shows the results of the in-distribution setting on the RAVEL task using Llama-3-8B, Llama-2-13b-chat-hf, and Llama-3-70B-Instruct.
@@ C.2 Ravel OOD with Two Shots
First N Token $T$=1 $0.903_{\hskip 1.0pt\pm 0.011}$ $0.860_{\hskip 1.0pt\pm 0.011}$ $0.893_{\hskip 1.0pt\pm 0.008}$ $0.876_{\hskip 1.0pt\pm 0.009}$ $0.866_{\hskip 1.0pt\pm 0.005}$ $0.789_{\hskip 1.0pt\pm 0.012}$ $0.693_{\hskip 1.0pt\pm 0.012}$
$T$=2 $\mathbf{0.933}_{\hskip 1.0pt\pm 0.004}$ $0.902_{\hskip 1.0pt\pm 0.007}$ $0.911_{\hskip 1.0pt\pm 0.006}$ $0.911_{\hskip 1.0pt\pm 0.009}$ $0.898_{\hskip 1.0pt\pm 0.006}$ $0.792_{\hskip 1.0pt\pm 0.015}$ $0.739_{\hskip 1.0pt\pm 0.011}$
Answer Token $T$=1 $0.881_{\hskip 1.0pt\pm 0.013}$ $0.861_{\hskip 1.0pt\pm 0.008}$ $0.889_{\hskip 1.0pt\pm 0.010}$ $0.877_{\hskip 1.0pt\pm 0.009}$ $0.848_{\hskip 1.0pt\pm 0.004}$ $0.746_{\hskip 1.0pt\pm 0.018}$ $0.651_{\hskip 1.0pt\pm 0.013}$
$T$=2 $0.889_{\hskip 1.0pt\pm 0.004}$ $0.898_{\hskip 1.0pt\pm 0.005}$ $0.894_{\hskip 1.0pt\pm 0.003}$ $0.901_{\hskip 1.0pt\pm 0.013}$ $0.827_{\hskip 1.0pt\pm 0.004}$ $0.673_{\hskip 1.0pt\pm 0.029}$ $0.504_{\hskip 1.0pt\pm 0.014}$

### v2

@@ 1 Introduction
As language models are increasingly deployed in high-stakes, open-ended scenarios, it is critical that we have a handle on how to prevent incorrect or unsafe outputs across a wide range of unanticipated user inputs. However, behavioral testing usually cannot be comprehensive given the compositionally large input space of a natural language task and the potential for distribution shifts across deployed environments. The traditional approach is to estimate the correctness of model predictions using confidence scores (Hendrycks & Gimpel, 2017; Lakshminarayanan et al., 2017; Guo et al., 2017). A more recent line of work uses internal representations of the model (usually supervised probes or use other statistical measures of model activations) to predict the model’s correctness on factual retrieval tasks (Azaria & Mitchell, 2023; Zou et al., 2023; Orgad et al., 2025; Marks & Tegmark, 2024; Gottesman & Geva, 2024; Ferrando et al., 2025; Ji et al., 2024; Yin et al., 2024; Chuang et al., 2024).
@@ 2.2 Correctness Estimation
Correctness estimation is a fundamental problem in machine learning evaluation and reliability assessment. It is the problem of predicting whether a model output matches the ground truth. For a classifier, a natural choice is to use probabilistic predictions, i.e., the confidence scores (Hendrycks & Gimpel, 2017; Lakshminarayanan et al., 2017). However, LMs’ outputs are structured predictions which add complexities to calibration (Kuleshov & Liang, 2015), and deep neural models are only well-calibrated to in-domain inputs after temperature scaling (Guo et al., 2017; Desai & Durrett, 2020; Kadavath et al., 2022; OpenAI, 2024; Wang et al., 2024). Alternatives have been proposed, such as using verbalized confidence scores for instruction-tuned models (Kadavath et al., 2022; Tian et al., 2023). Most related to ours is the line of work using internal representations to estimate the knowledge stored in LMs (Gottesman & Geva, 2024; Marks & Tegmark, 2024; Ferrando et al., 2025), assess the risk of hallucination (Azaria & Mitchell, 2023; Zou et al., 2023; Du et al., 2024; Chuang et al., 2024; Ji et al., 2024; Yin et al., 2024; Orgad et al., 2025), and predict jailbreaking behaviors (Arditi et al., 2024). However, identifying robust features from internal representations that generalize across tasks and datasets remains a challenge (Orgad et al., 2025). This is the problem that we focus on in this work.
@@ 5 Experiments
Confidence Score $\times N$
Correctness Probing Maybe $\times 0.5$NN
Counterfactual Simulation
Localization
Only
$\times NK$
Value Probing
Localization
Only
Maybe $\times 0.5$NN
Table 1: A comparison of methods and their requirements. Requires training: whether the method requires learning additional parameters; requires wrong samples: whether supervision from incorrect examples is needed; requires counterfactuals: whether supervision from counterfactual examples is needed; requires decoding: whether decoding at inference time is needed; inference cost: expected number of forward passes needed, with NN decoding steps and $K$ counterfactual samples. Localization only indicates the condition is required only for localizing causal variables, but not for correctness prediction; Maybe indicates methods that can apply to both prompt tokens and decoded tokens.
@@
Confidence Score (First N Token) TT=1 $0.793_{\hskip 1.0pt\pm 0.011}$ $0.935_{\hskip 1.0pt\pm 0.004}$ $0.796_{\hskip 1.0pt\pm 0.023}$ $\mathbf{0.810}_{\hskip 1.0pt\pm 0.009}$ $0.682_{\hskip 1.0pt\pm 0.016}$
TT=2 $0.804_{\hskip 1.0pt\pm 0.010}$ $0.900_{\hskip 1.0pt\pm 0.016}$ $0.836_{\hskip 1.0pt\pm 0.011}$ $0.795_{\hskip 1.0pt\pm 0.002}$ $0.668_{\hskip 1.0pt\pm 0.023}$
Confidence Score (Answer Token) TT=1 $0.498_{\hskip 1.0pt\pm 0.002}$ $0.928_{\hskip 1.0pt\pm 0.006}$ $0.795_{\hskip 1.0pt\pm 0.018}$ $\mathbf{0.810}_{\hskip 1.0pt\pm 0.009}$ $0.664_{\hskip 1.0pt\pm 0.015}$
TT=2 $0.498_{\hskip 1.0pt\pm 0.002}$ $0.870_{\hskip 1.0pt\pm 0.016}$ $0.679_{\hskip 1.0pt\pm 0.015}$ $0.795_{\hskip 1.0pt\pm 0.002}$ $0.657_{\hskip 1.0pt\pm 0.016}$
@@
[…]
@@
[…]
TT=2 $0.498_{\hskip 1.0pt\pm 0.003}$ $0.714_{\hskip 1.0pt\pm 0.011}$ $0.752_{\hskip 1.0pt\pm 0.005}$ $0.558_{\hskip 1.0pt\pm 0.009}$ $0.873_{\hskip 1.0pt\pm 0.013}$ $0.651_{\hskip 1.0pt\pm 0.013}$ $\underline{0.761}_{\hskip 1.0pt\pm 0.017}$ $0.679_{\hskip 1.0pt\pm 0.009}$ $0.592_{\hskip 1.0pt\pm 0.000}$ $\underline{0.730}_{\hskip 1.0pt\pm 0.000}$
    (Answer Token) TT=1 $0.544_{\hskip 1.0pt\pm 0.012}$ $0.771_{\hskip 1.0pt\pm 0.020}$ $0.709_{\hskip 1.0pt\pm 0.017}$ $0.596_{\hskip 1.0pt\pm 0.007}$ $\underline{0.908}_{\hskip 1.0pt\pm 0.016}$ $39_{\hskip 1.0pt\pm 0.011} 0.707$$_{\hskip 1.0pt\pm 0.021} 0.770$$70_{\hskip 1.0pt\pm 0.028}$ $\mathbf{0.931}_{\hskip 1.0pt\pm 0.018}$ $0.491_{\hskip 1.0pt\pm 0.018}$
TT=2 $0.498_{\hskip 1.0pt\pm 0.003}$ $0.729_{\hskip 1.0pt\pm 0.016}$ $0.671_{\hskip 1.0pt\pm 0.024}$ $0.521_{\hskip 1.0pt\pm 0.004}$ $\underline{0.906}_{\hskip 1.0pt\pm 0.014}$ $0.504_{\hskip 1.0pt\pm 0.014}$ $\underline{0.761}_{\hskip 1.0pt\pm 0.017}$ $0.679_{\hskip 1.0pt\pm 0.009}$ $\underline{0.902}_{\hskip 1.0pt\pm 0.018}$ $0.484_{\hskip 1.0pt\pm 0.016}$
@@
    (Background Variables) $0.556_{\hskip 1.0pt\pm 0.003}$ $0.597_{\hskip 1.0pt\pm 0.008}$ $0.541_{\hskip 1.0pt\pm 0.006}$ $0.749_{\hskip 1.0pt\pm 0.017}$ $0.812_{\hskip 1.0pt\pm 0.057}$ $0.837_{\hskip 1.0pt\pm 0.007}$ $0.729_{\hskip 1.0pt\pm 0.013}$ $0.698_{\hskip 1.0pt\pm 0.020}$ $56_{\hskip 1.0pt\pm 0.009} 0.660$$60_{\hskip 1.0pt\pm 0.009}$
@@ 5.1 Setup
We construct a test set with an equal number of correctly and wrongly predicted inputs, which turns the correctness prediction task into a balanced binary classification task. Following the evaluation setup of similar truthfulness prediction tasks (Orgad et al., 2025; Yin et al., 2024), we measure the prediction accuracy using AUC-ROC.
@@ A.1 Task Prompts
Table LABEL:tab:task_prompts summarizes the templates and variables used to construct prompts. Each prompt consists of an instruction, a chat template (for chat models), and an optional prefix prepended to model outputs. Prompt design adheres to the following criteria:
@@
We select prompts that achieve reasonably high accuracy, but not perfect accuracy, as discussed in Section 3. This is meant to ensure that (1) the model uses some systematic solutions rather than relying on random guesses, and (2) the model makes some errors, enabling meaningful analysis of conditions under which the model makes correct or incorrect predictions (see task accuracy in Table 3). This balance is especially critical for simpler tasks like IOI, where today’s LLMs like Llama-3-8B-Instruct frequently achieve near-perfect performance.
@@
{upper_bound}>>{lower_bound} and that they are distinct from {query}.
@@
[{"市": "{city_a_in_japan}", "国": "Japan", {"市": "{city_b_not_in_japan}",
@@ B.2 Training Setup and Hyperparameters
Task IOI PriceTag RAVEL MMLU UnlearnHP
Intervention
Dimensions
256 1 1024 4 4
Intervention
Localizations
Variables
Position
$P$ and $Q$
Country Position Refusal
Layers
9-12
@@
13-14 17 18
@@
For counterfactual simulation, we use 10K pairs randomly sampled from 1024x1024 correct examples as the training data. We use the AdamW optimizer with constant LR = $10^{-4}$ with no weight decay, trained for one epoch.
@@
B.3 Choice of Variables in Counterfactual Simulation
We show the causal mechanisms and variables used in counterfactual simulation in Figure 6.
@@ C.1 Model Variations
             TT=1 $0.867_{\hskip 1.0pt\pm 0.019}$ $0.638_{\hskip 1.0pt\pm 0.036}$ $0.800_{\hskip 1.0pt\pm 0.016}$
             TT=2 $0.861_{\hskip 1.0pt\pm 0.023}$ $0.649_{\hskip 1.0pt\pm 0.020}$ $0.898_{\hskip 1.0pt\pm 0.013}$
@@
             TT=1 $0.853_{\hskip 1.0pt\pm 0.022}$ $0.593_{\hskip 1.0pt\pm 0.012}$ $0.766_{\hskip 1.0pt\pm 0.016}$
             TT=2 $0.724_{\hskip 1.0pt\pm 0.035}$ $0.607_{\hskip 1.0pt\pm 0.014}$ $0.769_{\hskip 1.0pt\pm 0.019}$
@@
Table 6 shows the results of the in-distribution setting on the RAVEL task using Llama-3-8B, Llama-2-13b-chat-hf, and Llama-3-70B-Instruct. Overall, we do not see scaling up model size significantly improves the confidence scores baseline.
@@ C.2 Ravel OOD with Two Shots
First N Token TT=1 $0.903_{\hskip 1.0pt\pm 0.011}$ $0.860_{\hskip 1.0pt\pm 0.011}$ $0.893_{\hskip 1.0pt\pm 0.008}$ $0.876_{\hskip 1.0pt\pm 0.009}$ $0.866_{\hskip 1.0pt\pm 0.005}$ $0.789_{\hskip 1.0pt\pm 0.012}$ $0.693_{\hskip 1.0pt\pm 0.012}$
TT=2 $\mathbf{0.933}_{\hskip 1.0pt\pm 0.004}$ $0.902_{\hskip 1.0pt\pm 0.007}$ $0.911_{\hskip 1.0pt\pm 0.006}$ $0.911_{\hskip 1.0pt\pm 0.009}$ $0.898_{\hskip 1.0pt\pm 0.006}$ $0.792_{\hskip 1.0pt\pm 0.015}$ $0.739_{\hskip 1.0pt\pm 0.011}$
Answer Token TT=1 $0.881_{\hskip 1.0pt\pm 0.013}$ $0.861_{\hskip 1.0pt\pm 0.008}$ $0.889_{\hskip 1.0pt\pm 0.010}$ $0.877_{\hskip 1.0pt\pm 0.009}$ $0.848_{\hskip 1.0pt\pm 0.004}$ $0.746_{\hskip 1.0pt\pm 0.018}$ $0.651_{\hskip 1.0pt\pm 0.013}$
TT=2 $0.889_{\hskip 1.0pt\pm 0.004}$ $0.898_{\hskip 1.0pt\pm 0.005}$ $0.894_{\hskip 1.0pt\pm 0.003}$ $0.901_{\hskip 1.0pt\pm 0.013}$ $0.827_{\hskip 1.0pt\pm 0.004}$ $0.673_{\hskip 1.0pt\pm 0.029}$ $0.504_{\hskip 1.0pt\pm 0.014}$
@@
Appendix D Limitations
Generalizability to non-linear representations
Requirement of Prior Knowledge of the Task
Indeed, many real-world applications already exhibit shared structures readily exploitable by our methods. Two practical domains exemplify this. The first is LLM evaluation: constructing benchmarks typically requires costly human annotations; thus, a reliable correctness estimator helps benchmark creators prioritize uncertain or error-prone examples. Meanwhile, systematic evaluation often requires structured input—for example, SQuAD adversarial (Jia & Liang, 2017) or GSM-symbolic (Mirzadeh et al., 2025)—which inherently reflects high-level task structures readily exploitable by our framework. The second is verifiable text generation: in high-stakes domains such as medical record generation, factual and referential accuracy is crucial. One strategy is having LLMs generate symbolic references to structured source data that are easy to verify (Hennigen et al., 2024). Such symbolic references act as templated prompts convertible into high-level models, enabling our methods to estimate the correctness of texts produced by these decoding algorithms.